# Family Unification in $SO(16)$ Grand Unification

### v1

@@ 7 Renormalization group equation for gauge coupling constant
where Vector, Weyl, and Real under the summation symbol stand respectively for real vector, Weyl fermion, and real scalar fields in terms of 4D theories. The summation is done for particles with mass smaller than the specified energy scale $\mu$. The vector bosons are the gauge bosons, and thus belong to the adjoint representation 120 of $SO(16)$ with $T(R_{V}={\bf 120})=C_{2}(SO(16))=14$, where $C_{2}(SO(16))$ is the quadratic Casimir invariant of the adjoint representation of $SO(16)$, and $T(R_{X})$ is the Dynkin index of the irreducible representation $R_{X}$ of $SO(16)$. Finally, $T({\bf 128})=16$ [38].
Let us now consider the RGE for the 4D gauge coupling constant in 6D gauge theories given in Eq. (7.1) by using the $\beta$ function coefficient given in Eq. (7.2), which depends on its field content at the energy scale $\mu$. For the $\beta$ function coefficient, we take into account the contribution from not only the zero modes but also the KK modes with masses less than the renormalization scale $\mu$. Once specifying the mass spectrum of the model, we can calculate the RGE for the gauge coupling constant at one-loop level. In general, it is difficult to write down the exact mass spectrum because it depends on orbifold boundary conditions as well as the parameters of 6D bulk and 4D localized terms. At the leading-order approximation, the masses of zero modes and 4D fields are much less than $\mu$ and those of the $k$-th KK modes can be set as $m=km_{KK}$. By using this mass spectrum, the RGE of the gauge coupling constant can be divided into two regimes:
@@ 8 Summary and discussions
In this paper, we have proposed a unified model for the SM gauge symmetry and family symmetry based on the $SO(16)$ gauge symmetry on the 6D spacetime $M^{4}\times D^{2}/\mathbb{Z}_{N}$ $(N\geq 10,N\in\mathbb{Z}_{\rm even})$. In this model, the three chiral generations of quarks and leptons in the SM can be unified into one 6D Weyl fermion in the spinor representation of $SO(16)$ when considering the symmetry breaking pattern $SO(16)\to SO(10)\times SU(3)\times U(1)$. The 6D $SO(16)$ gauge anomaly is canceled due to the vectorlike nature of the Weyl fermion in the model, and the 4D gauge anomalies are canceled by introducing suitable 4D localized fermions at the fixed point. The gauge coupling constant of $SO(16)$ is found to be asymptotically free above the KK scale.
@@
Finally, we make a brief comment on the masses of quarks and leptons. The analysis of mass matrices in models with $SU(3)$ family symmetry has already been done in Refs. [84, 85, 86, 87, 88, 89, 6, 90, 91, 92, 93, 94, 95]. Though phenomenologically important, a detailed study of masses and mixing in this model, however, is outside the scope of this paper and will be discussed in a future work.

### v2

@@ 7 Renormalization group equation for gauge coupling constant
where Vector, Weyl, and Real under the summation symbol stand respectively for real vector, Weyl fermion, and real scalar fields in terms of 4D theories. The summation is done for particles with mass smaller than the specified energy scale $\mu$. The vector bosons are the gauge bosons and thus belong to the adjoint representation 120 of $SO(16)$ with $T(R_{V}={\bf 120})=C_{2}(SO(16))=14$, where $C_{2}(SO(16))$ is the quadratic Casimir invariant of the adjoint representation of $SO(16)$, and $T(R_{X})$ is the Dynkin index of the irreducible representation $R_{X}$ of $SO(16)$. Finally, $T({\bf 128})=16$ [39].
Let us now consider the RGE for the 4D gauge coupling constant in the 6D gauge theories given in Eq. (7.1) by using the $\beta$ function coefficient given in Eq. (7.2), which depends on its field content at the energy scale $\mu$. For the $\beta$ function coefficient, we take into account the contribution from not only the zero modes but also the KK modes with masses less than the renormalization scale $\mu$. Once specifying the mass spectrum of the model, we can calculate the RGE for the gauge coupling constant at one-loop level. In general, it is difficult to write down the exact mass spectrum because it depends on orbifold boundary conditions as well as the parameters of 6D bulk and 4D localized terms. At the leading-order approximation, the masses of zero modes and 4D fields are much less than $\mu$ and those of the $k$-th KK modes can be set as $m=km_{KK}$. By using this mass spectrum, the RGE of the gauge coupling constant can be divided into two regimes:
@@ 8 Summary and discussions
In this paper, we have proposed a unified model for the SM gauge symmetry and family symmetry based on the $SO(16)$ gauge symmetry on the 6D spacetime $M^{4}\times D^{2}/\mathbb{Z}_{N}$ $(N\geq 9)$. In this model, the three chiral generations of quarks and leptons in the SM can be unified into one 6D Weyl fermion in the spinor representation of $SO(16)$ when considering the symmetry-breaking pattern $SO(16)\to SO(10)\times SU(3)\times U(1)$. The 6D $SO(16)$ gauge anomaly is canceled due to the vectorlike nature of the Weyl fermion in the model, and the 4D gauge anomalies are canceled by introducing suitable 4D localized fermions at the fixed point. The gauge coupling constant of $SO(16)$ is found to be asymptotically free above the KK scale.
@@
Finally, we make a brief comment on the masses of quarks and leptons. The analysis of mass matrices in models with $SU(3)$ family symmetry has already been done in Refs. [85, 86, 87, 88, 89, 90, 6, 91, 92, 93, 94, 95, 96]. Though phenomenologically important, a detailed study of masses and mixing in this model, however, is outside the scope of this paper and will be discussed in future work.